For the reaction 2 A(g) + 3 B(g) <-->2 C(g) + 2 D(g), the equilibrium constant at 298 K is 1.605×10-11. If the initial partial pressures of A, and B are 0.564 bar, and 0.728 bar, respectively, determine the partial pressure of D (in bar) once equilibrium has been reached at 298 K.
Answer #1

K=1-60541011 2A + 38 = 2C + RD, - initial 0.564 bar 0.728 baso o final (0.564-2x) (0=728-32e) . 2e are p = p- be? (22] (2232